Problem 8-18A Comprehensive Variance Analysis [LO8-4, LO8-5, LO8-6]

Miller Toy Company manufactures a plastic swimming pool at its Westwood Plant. The plant has been experiencing problems as shown by its June contribution format income statement below:

Budgeted Actual
  Sales (3,000 pools) $ 210,000    $ 210,000   
    
  Variable expenses:      
     Variable cost of goods sold* 38,220    49,235   
     Variable selling expenses 15,000    15,000   
    
  Total variable expenses 53,220    64,235   
    
  Contribution margin 156,780    145,765   
    
  Fixed expenses:      
     Manufacturing overhead 66,000    66,000   
     Selling and administrative 81,000    81,000   
    
  Total fixed expenses 147,000    147,000   
    
  Net operating income (loss)    $ 9,780    $ (1,235)
    
*Contains direct materials, direct labor, and variable manufacturing overhead.

Janet Dunn, who has just been appointed general manager of the Westwood Plant, has been given instructions to “get things under control.” Upon reviewing the plant’s income statement, Ms. Dunn has concluded that the major problem lies in the variable cost of goods sold. She has been provided with the following standard cost per swimming pool:

Standard Quantity or Hours Standard Price
or Rate
Standard Cost
  Direct materials    3.1 pounds $ 2.60 per pound $ 8.06   
  Direct labor    0.5 hours $ 7.20 per hour    3.60   
  Variable manufacturing overhead    0.4 hours* $ 2.70 per hour    1.08   
    
  Total standard cost $ 12.74   
    
*Based on machine-hours.
     During June the plant produced 3,000 pools and incurred the following costs:
a.

Purchased 14,300 pounds of materials at a cost of $3.05 per pound.

b.

Used 9,100 pounds of materials in production. (Finished goods and work in process inventories are insignificant and can be ignored.)

c. Worked 2,100 direct labor-hours at a cost of $6.90 per hour.
d.

Incurred variable manufacturing overhead cost totaling $4,650 for the month. A total of 1,500 machine-hours was recorded.

It is the company’s policy to close all variances to cost of goods sold on a monthly basis.
Required:
1. Compute the following variances for June:
a.

Materials price and quantity variances. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance).)

            

b.

Labor rate and efficiency variances. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance).)

            

c.

Variable overhead rate and efficiency variances. (Do not round your intermediate calculations. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance).)

            

2.

Summarize the variances that you computed in (1) above by showing the net overall favorable or unfavorable variance for the month. (Input all values as positive amounts. Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance).)

        

3.

Pick out the two most significant variances that you computed in (1) above. (You may select more than one answer. Single click the box with a check mark for correct answers and double click to empty the box for the wrong answers.)

Materials price variance
Labor efficiency variance
Variable overhead efficiency variance
Labor rate variance
Variable overhead rate variance
Materials quantity variance
Answer #1

1. variance for June

(a)

Material price variance = AQ (AP - SP)

= 14300 pounds (3.05 - 2.60)

Material price variance = 6435 U

Material quantity variance = SP (AQ - SQ)

= 2.60 (9100 - 9300*)

Material quantity variance = 520 F

*3000 pools x 3.1 per pound = 9300

(b) Labor rate variance = AH (AR - SR)

= 2100 pounds (6.90 - 7.20)

Labor rate variance = 630 F

Labor efficiency variance = SR (AH - SH)

= 9.20 (2100 - 1500*)

Labor efficiency variance = 4320 U

*3000 pools x 0.5 per pound = 1500

(c)  Variable overhead rate variance = AH (AR - SR)

= 1500 (3.1* - 2.70)

Variable overhead rate variance = 600 F

* 4650/1500 hours = 3.1 per hour

Variable overhead efficiency variance = AH (AR - SR)

= 2.70 (1500 - 1200*)

Variable overhead efficiency variance = 810 U

*3000 pools x 0.4 per pound = 1200

2. summarising the variance

Material price variance = 6435 U

Material quantity variance = 520 F

Labor rate variance = 630 F

Labor efficiency variance = 4320 U

Variable overhead rate variance = 600 F

Variable overhead efficiency variance = 810 U

Net variance for the month = $ 9815U  

This will reduce the profit in the income statement

3. Two most significant variance

Material price variance = 6435 U

Labor efficiency variance = 4320 U

Change in transportation expenses can change the material available for use and if the labor efficiency has become worst which lead to unfavorable labor efficiency variance.